OpenStudy (anonymous):

If y = 61 when x = –305, what is y when x = –315?

OpenStudy (anonymous):

The question is in the wrong subject -_- Answer is 63 anyways

OpenStudy (anonymous):

@deontrell \[\frac{ 61 }{ -305}=\frac{ y }{ -315 }\] Cross multiply and you'll get: -19, 215 = -305y Divide both sides by (-305): \[\frac{ -19, 215 }{ -305 }=\frac{ -305y }{-305 }\] Use cancellation: \[y=63\] Hope this helps! :)

OpenStudy (swag):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. If you need any help feel free to message me or any other ambassador. Welcome to the OpenStudy Community! Ambassador SWAG